# Randomly sampling bipartite networks with fixed degree sequences

### v3

@@ 5 Discussion
Researchers performing statistical analysis of a bipartite network need to draw a random sample of all bipartite networks with the same degree sequences as the observed network. Efficient ‘trade’ algorithms exist for sampling bipartite networks through repeated randomization trades [Strona et al., 2014, Godard & Neal, 2022]. However, it is difficult to know how many trades needed to ensure the sample is random. In this paper, I proposed a stopping rule that focuses on the distribution of the sampled networks’ distances from a starting network, and stops performing trades once this distribution stabilizes. For a wide range of degree sequences, this stopping rule ensures with high probability (93.2% on average) that the sample is a random sample of all bipartite networks with the given degree sequence. Additionally, in three larger empirical examples, the stopping rule performs as expected, and returns a sample in a practical amount of time.
@@
The stopping rule offers a solution to a practical challenge encountered by researchers: determining when a sample of bipartite networks is random, regardless of how many trades this might require. These analyses do not directly answer a related question: how many trades are required, or what is the mixing time of a trade algorithm? While the mixing time of trade algorithms remains unknown, these results offer some insight, while the stopping rule offers a tool for future research on their mixing time. Consistent with prior work [Strona et al., 2014, Carstens, 2015], the number of trades required to achieve a random sample according to the stopping rule is closely associated with the number of top-nodes ($r=0.979$). However, whereas prior work has suggested five times the number of top-nodes as an approximate mixing time [Strona et al., 2014], these results suggest this is an underestimate. This rule-of-thumb suggests that 25 trades should be sufficient for a network containing with five top-nodes, however the stopping rule required an average of 31.6 trades. Similarly, the rule-of-thumb suggests that 500 trades should be sufficient for the Senate network containing 100 top-nodes, however the stopping rule required 1400 trades. While prior approximations of trade algorithms’ mixing time may be too low, the stopping rule may provide a way for future research to estimate mixing time more accurately and for larger degree sequences.
These results must be interpreted in light of three limitations. First, it was only possible to evaluate the performance of the stopping rule for degree sequences that define a $\mathcal{B}$ for which $|\mathcal{B}|\leq 100$. This is primarily a practical computational limitation; with additional computing time, it is theoretically possible to evaluate it for longer degree sequences. However, $|\mathcal{B}|$ can grow very rapidly as a function of network size and in most cases is unknown [Barvinok, 2010], making it difficult to enumerate all members of $\mathcal{B}$.
@@
Second, although the stopping rule provides confidence that the resulting sample is random, this confidence comes at the cost of greater time and memory requirements. In terms of time, the additional computations necessary to use the stopping rule (i.e., periodically computing $D$ and a KS test) take approximately 3.5 times longer than the trade algorithm alone. In terms of memory, if the required number of trades was known in advance, then a trade algorithm can be used to sample networks from $\mathcal{B}$ one at a time. In contrast, using the stopping rule requires holding all the sampled networks in memory as they are incrementally randomized and compared in terms of their distance from the starting network. However, as noted above, future research may use the stopping rule to investigate the mixing time of trade algorithms, thereby providing guidance on choosing a number of trades in advance and eliminating the need for the stopping rule.
Finally, although on average the stopping rule yields a random sample in 93.2% of cases, this still means that 6.8% of the time it yields a non-random sample. While the risk that the stopping rule yields a non-random sample should not be ignored, it is worth noting that the sample is likely still nearly-random and that this level of error is close to the error tolerance (e.g., $\alpha=0.05$) typically accepted in conventional statistical inference. When greater certainty of a sample’s randomness is necessary, one can impose a higher threshold on regarding the KS test as non-significant. The results presented in table 1 view the KS test used by the stopping rule as non-significant when $p>0.05$. However, if a $p>0.95$ threshold is used instead, then on average the stopping rule yields a random sample in 95.17% of cases. However, this increased certainty comes at a computational cost of requiring more trades (on average 33 trades vs. 23).
@@
All data and code necessary to replicate this study are available at https://osf.io/rmwpz/.

### v4

@@ 5 Discussion
Researchers performing statistical analysis of a bipartite network need to draw a random sample of all bipartite networks with the same degree sequences as the observed network. Efficient ‘trade’ algorithms exist for sampling bipartite networks through repeated randomization trades [Strona et al., 2014, Godard & Neal, 2022]. However, it is difficult to know how many trades needed to ensure the sample is random. In this paper, I proposed a stopping rule that focuses on the distribution of the sampled networks’ distances from a starting network, and stops performing trades once this distribution stabilizes. For a wide range of degree sequences, this stopping rule ensures with high probability (87.5% on average, 92% adjusted for type-I error) that the sample is a random sample of all bipartite networks with the given degree sequence. Additionally, in three larger empirical examples, the stopping rule performs as expected, and returns a sample in a practical amount of time.
@@
The stopping rule offers a solution to a practical challenge encountered by researchers: determining when a sample of bipartite networks is random, regardless of how many trades this might require. These analyses do not directly answer a related question – how many trades are required, or what is the mixing time of a trade algorithm – but they do offer some insight. In their provided code, Strona et al. [2014] implicitly suggested that $5T$ trades was sufficient, where $T$ is the number of top nodes, but they did not provide any rationale or evidence for this suggestion. Was their suggestion reasonable?
For the small networks described in section 4.1, an average of about 4 trades per top node were required to obtain a random sample ($M=4.11,SD=0.50$). These results suggest that a $5T$ rule-of-thumb may be slightly liberal, but not unreasonable, in small networks. In contrast, for the empirical networks described in section 4.2, many more trades were required: $9T$ for Darwin’s Finches, $7T$ for Southern Women, and $11T$ for the US Senate. These results suggests that in larger networks a $5T$ rule-of-thumb may be too conservative, and that more trades are needed to achieve a random sample. While the results presented above do not provide definitive guidance on how many trades to perform in general, the stopping rule provides a way to determine how many trades to perform in a specific case, as well as a tool for further research on trade algorithm mixing times.
@@
These findings must be interpreted in light of two computational limitations. First, because $|\mathcal{B}|$ grows very rapidly as a function of network size and in most cases is unknown [Barvinok, 2010], it is only possible to evaluate the randomness of a sample for relatively small bipartite networks. Therefore, although the results presented in section 4.2 are consistent with the resulting samples being random, their randomness cannot be evaluated directly. Second, although the stopping rule provides confidence that the resulting sample is random, this confidence comes at the cost of greater time and memory requirements. In terms of time, using the stopping rule takes longer than simply performing trades because it requires periodically computing $D$ and performing a AD or DTS test. In terms of memory, using the stopping rule requires requires holding all the sampled networks in memory as they are incrementally randomized and compared in terms of their distance from the starting network. However, future research may use the stopping rule to investigate the mixing time of trade algorithms, thereby providing guidance on choosing a number of trades in advance and eliminating the need for the stopping rule itself.
@@
Supplementary information, and all data and code necessary to replicate these results, are available at https://osf.io/rmwpz/?view_only=5d18c67044b44474addba3824cab2aa5.